OpenStudy (anonymous):

A line has a slope of 7 and contains the point (5, −9). Which equations represent the line? Choose all answers that are correct

OpenStudy (anonymous):

A. −7x + y = −44 B. 7x – y = 44 C. (y − 9) = 7(x + 5) D. (y + 9) = 7(x − 5)

OpenStudy (jojo4eva):

the answer is D

OpenStudy (anonymous):

thanks

OpenStudy (jojo4eva):

wait wait there might be more

OpenStudy (jojo4eva):

give me a sec

OpenStudy (anonymous):

@popcorn23 I know B. is one. I thought A. was one to but i could be wrong. @jojo4eva Im sure the smart one could verify it for ya. -_-

OpenStudy (anonymous):

lol ok thanks @keke_luvs_u

OpenStudy (anonymous):

@popcorn23 Lol im not smart. Welcome (:

OpenStudy (jojo4eva):

yup shes right @popcorn23

OpenStudy (anonymous):

y=7x-44

OpenStudy (anonymous):

lol see you are smart @keke_luvs_u

OpenStudy (anonymous):

@popcorn23 It was luck but k.

OpenStudy (jojo4eva):

Actually D is the only right answer ir seems

OpenStudy (jojo4eva):

No A is right also. So A and D are your answers

OpenStudy (anonymous):

ok thanks lol ive never had people actually help me :) @keke_luvs_u @jojo4eva

OpenStudy (jojo4eva):

no problem if you need anything else dont hesitate to tag me. And yup thats what we are here for

OpenStudy (anonymous):

@popcorn23 Welcome haha (;